1. If persons not in the labor force total 4 million and persons in the civilian labor force total 60 million, then how many persons are unemployed?

a.57 million

b.64 million

c.20 million

d.There is not enough information to answer the question.

( for this question I pick b).

2.

Exhibit 6-1

Year

CPI

1

107.6

2

106.6

3

113.6

4

118.3

Refer to Exhibit 6-1. Prices changed by _________ percent between year 1 and year 2.

a.1.0

b.-0.9

c.-2.1

d.-1.6

( for this I pick A)

3.

The CPI was 213 at the beginning of year 1 and 234 at the end of year 1. By how much did prices rise during the year?

a.10.12 percent

b.9.86 percent

c.11.34 percent

d.8.10 percent

e.none of the above

( for this ques. i pick B.)

4. Suppose that the frictional unemployment rate is 2 percent, the natural unemployment rate is 3.5 percent, and the actual unemployment rate is 7.5 percent. The structural unemployment rate is ____________ and the cyclical unemployment rate is __________.

a.5.5 percent, 2 percent

b.1.5 percent, 4 percent

c.4 percent, 9.5 percent

d.1.5 percent, 5.5 percent

( I chose option b)

5. In an economy with a population of 70 million persons, 35 million are employed as civilians and 3 million are not employed but are both actively looking for a job and available for work. The unemployment rate in this economy is approximately

a.7.9%.

b.8.6%.

c.4.3%.

d.10.1.

6. Suppose that an economy has a civilian noninstitutional population of 50,000 persons. There are 30,000 persons in the civilian labor force, of which 25,000 are employed. How many persons in this economy are unemployed?

a.5,000

b.20,000

c.25,000

d.35,000

( for this i pick A)

7.

The frictional unemployment rate is 3 percent, the structural unemployment rate is 3.3 percent, and the (actual) unemployment rate is 6 percent. It follows that the natural unemployment rate is ___________ percent and the cyclical unemployment rate is ____________ percent.

a.-0.3 percent, 6.3 percent

b.9 percent, -0.3 percent

c.6.3 percent, -0.3 percent

d.0.3 percent, 6.3 percent

e.none of the above

( for this I pick option C)
Answer #1

1. d.There is not enough information to answer the question.
(We need information about how many people are employed.)

2. b.-0.9
(Percentage change in prices = [(106.6-107.6)/107.6]*100 = -100/107.6 = -0.9%)

3. b.9.86 percent
(Percentage change in prices = [(234-213)/213]*100 = (21/213)*100 = 9.86%)

4. b.1.5 percent, 4 percent
(Structural unemployment rate = natural unemployment rate - frictional unemployment rate = 3.5 - 2 = 1.5%
Cyclical unemployment rate = actual unemployment rate - natural unemployment rate = 7.5 - 3.5 = 4%)

5. a.7.9%
(Unemployment rate = [3/(35+3)]*100 = (3/38)*100 = 7.9%)

6. a.5,000
(Unemployed = Labor force - Employed = 30,000 - 25,000 = 5,000)

7. c.6.3 percent, -0.3 percent
(Natural unemployment rate = Structural unemployment rate + frictional unemployment rate = 3 + 3.3 = 6.3%
Cyclical unemployment rate = Actual unemployment rate - Natural unemployment rate = 6 - 6.3% = -0.3%)